OpenStudy (anonymous):

Hello Guys, can someone help me to understand why the integral of e^-x is -e^-x and vice versa. Thanks for your help.

myininaya (myininaya):

\[\int\limits_{}^{}e^{-x} dx\] let u=-x => du=-dx so we have \[\int\limits_{}^{}e^{u} (-du) =- \int\limits_{}^{}e^u du=- e^u+C=-e^{-x}+C\]

OpenStudy (anonymous):

If you want proof put e^-x itslef as t x=-ln t dx=-(1/t)dt so t and t cancel you have just inegral of -dt which is -t=-e^-x

myininaya (myininaya):

its hard for me to read that

OpenStudy (anonymous):

Thank you very much. I was confuse between e^x and e^-x. I know that integral of e^x is e^x, and was thinking if integral of e^-x is the same as e^-x. But i know now that if it is not e^x then i must do substitution. I hope my comments are correct.

myininaya (myininaya):

So you look like you made substitution

myininaya (myininaya):

\[x=-\ln(t) \]

myininaya (myininaya):

\[dx=-\frac{1}{t} dt\]

OpenStudy (anonymous):

@Myininaya i didn't do the substitution. i just wrote what its in the book. I didn't understand why the author got that answer. But your answer made it clear to me

OpenStudy (anonymous):

@myininaya are u telling me? yeah i did make that substitution, if you wanted proof that integral of e^x is e^x only i'm not too good with Writing those equations visually...

myininaya (myininaya):

\[\int\limits_{}^{}e^{-(-\ln(t))} \frac{-1}{t} dt=-\int\limits_{}^{}e^{\ln(t)} \frac{1}{t} dt=- \int\limits_{}^{}t \cdot \frac{1}{t} dt=- \int\limits_{}^{} 1 dt =- t+c\]

myininaya (myininaya):

so if \[x=-\ln(t) => -x=\ln(t)=>e^{-x}=e^{\ln(t)}=>e^{-x}=t\]

myininaya (myininaya):

so both ways work

myininaya (myininaya):

\[-e^{-x}+c\]

myininaya (myininaya):

gj i like the first way that had above better though

OpenStudy (anonymous):

err... whatever... \m/

myininaya (myininaya):

shank i wasn't saying you were wrong i was just having trouble reading it i wanted to see it clearly sorry

OpenStudy (anonymous):

I never said you said i was wrong :) \m/

myininaya (myininaya):

ok :)

myininaya (myininaya):

you don't have to make substitutions once you get use to what these answers look like

OpenStudy (anonymous):

Yeah of course i just thought he/she wanted the proof :D

myininaya (myininaya):

right she has two proofs now your proof and my proof

myininaya (myininaya):

she/he*